How much heat is released when a mixture of 3.34 g B2H6 and 2.22 g O2 is burned?
1
Expert's answer
2019-04-15T07:55:14-0400

"B2H6(g)+3O2(g) \to B2O3(s)+3H2O(g)–2035kJ"

"\dfrac{3.34g}{27.67g/mol}=0.1207moles(B2H6)"

"\dfrac{2.22g}{32g/mol}=0.06938moles(O2)"

"\dfrac{1mole(B2H6)*0.06938moles(O2)}{3moles(O2)}=0.0231moles(B2H6)"

"0.0231moles(B2H6)*2035kJ\/mol=47kJ"